# Steady-state phase transition in one-dimensional quantum contact process

### v2

@@ Abstract
However, when the model is extended to the quantum realm, the new features arise due to the interplay between quantum coherent time-evolution and dissipation. Early investigations into the quantum contact process (QCP) focused on its non-equilibrium phase transition and its distinctions from the classical directed percolation class Marcuzzi2016 ; Buchhold2017 ; Jo2019 ; Roscher2018 ; Carollo2019 ; Gillman2019 ; Gillman2020 ; Jo2021 . A remarkable feature of the QCP is that it can show phase transition in one-dimensional system, however the nature of such a phase transition remains controversial Carollo2019 ; Gillman2019 ; Marcuzzi2016 ; Buchhold2017 ; Jo2019 ; Roscher2018 . Whether the phase transition is continuous or not still remains a matter of debate, and constitutes the main focus of the present work.
@@
Apart from the Liouvillian gap $\mu_{0}$, the structure of the Liouvillian eigenvalues also reflects the time-scale of relaxation towards the steady state. This is quite relevant to the consuming time in a practical numerical simulation. In Fig. 2(b), it is shown the first few largest negative real parts of Liouvillian eigenvalues for $\Omega/\Gamma=5.8$ which is located near the transition point. One can see that the largest negative real part (blue circle) approaches zero as $1/L\rightarrow 0$, as indicated by the arrow Fig. 2(a). Moreover the second largest negative real part (red circles) is constant 0.5-0.5 and it bounds the real parts of all the rest eigenvalues. Namely, the eigenvalues with zero and the largest negative real parts are isolated from the other eigenvalues when the system size becomes larger and larger, signaling a metastability during the time-evolution Macieszczak2016 ; Rose2016 ; Macieszczak2021 .
In the inset of Fig. 2(b), it is shown that the time-evolution of the averaged population $\langle\bar{n}(t)\rangle$ for $\Omega/\Gamma=5.03$ and $5.035$. Starting with the full active state $|\uparrow\uparrow$$\uparrow\rangle$, the $\langle\bar{n}(t)\rangle$ in both cases decreases fast in the early stage. However it stays in a plateau with finite population for a pretty long time ($10^{2}\Gamma^{-1}$) and then dramatically falls onto the eventual absorbing state. The existence of the metastable state, which is not observed in the (single-site) mean-field approximation, is a consequence of the inclusion of short-range correlations in the system. It reminds us that one should prolong the numerical simulation time to ensure that the system escapes from the metastable state.
@@
Figure 3: (a) The $\langle\hat{n}_{1}\rangle_{\text{ss}}$ as a function of the effective field $F^{n}$ for the CMF simulation with $L=7$. The slope of the thin solid line is one. The intersection of the thin solid line and the curve of $\langle n_{1}\rangle_{\text{ss}}$ denotes the steady-state solution. The intersection denoted by full (empty) magenta diamond indicates the stable (unstable) steady-state solution. (b) The time-evolution trajectories of the reduced state of the boundary site originating from different initial states $|\psi\rangle=\bigotimes_{j}{|\psi\rangle_{j}}$ where the identical $|\psi_{j}\rangles$ locate at the surface of the Bloch sphere. The diamonds and the circle correspond to the intersections in (a). (c) The branches of $\langle\bar{n}\rangle_{\text{ss}}$ as a function of $\Omega/\Gamma$ in the CMF approximation with $L=7,9$ and $11$. The upper solid (lower dashed) curves denote the stable (unstable) active states. The red solid line denotes the absorbing state. (d) The resulting CMF transition point $\Omega_{c}(L)/\Gamma$ for different $L$. The green and blue symbols denote the odd- and even-size clusters, respectively. The square in magenta marks the transition point $\Omega_{c}/\Gamma=5.83$ through the Liouvillian gap analysis in Fig. 2(a).
@@
Fig. 3(a) shows the steady-state population of the boundary site $\langle\hat{n}_{1}\rangle_{\text{ss}}$ as a function of $F^{n}$ for $L=7$. One can see that for $\Omega/\Gamma=4$ the curve of $\langle\hat{n}\rangle_{\text{ss}}$ only intersects with the straight line $F^{n}=\langle\hat{n}_{1}\rangle_{\text{ss}}$ at the origin meaning that the unique steady state is the absorbing state. While for $\Omega/\Gamma=6$ there are two more intersections at $F^{n}=0.0725$ and $0.3085$ marked by the magenta diamonds. The former (with the derivative greater than one) is an unstable fixed point while the latter indicates the physical steady state. As shown in Fig. 3(b), the unstable steady state is brought to the absorbing state by a random perturbation while stable steady states tends to attract all the time-evolution trajectories originating from different initial states on the surface of the Bloch sphere. The stable and unstable steady-state solutions for different $L$ are shown in Fig. 3(c). The feature of the saddle-node bifurcation of the steady-state population is preserved in the CMF results. In particular, the saddle-node point $S_{2}$ remains infinity, while the $S_{1}$ shifts right towards to the transition point in TDL, which evidences the discontinuous phase transition.
The CMF transition points $\Omega_{c}(L)$ for different sizes of cluster is shown in Fig. 3(d). Due to the strong boundary effect in small clusters, there is an even-odd effect in the behavior of the $\Omega_{c}(L)$ as a function of $L$. For odd-size clusters, the $\Omega_{c}(L)$ increases as the $L$ increasing while it exhibits conversely for the even-size clusters. However for larger $L$ the $\Omega_{c}(L)$ is expected to converge toward the transition point $\Omega_{c}/\Gamma\approx 5.83$ predicted by the Liouvillian spectrum analysis.
@@ Acknowledgements.
We warmly thank useful discussions with Prof. Fernando Iemini and Dr. Wen-Bin He. This work was supported by Natural Science Foundation of Liaoning Province No. 2025-MS-009 and the Natural Science Foundation of China under Grants No. 11975064. X. L. is supported by the Hong Kong RGC Early Career Scheme (Grant No. 24308323), Guangdong Provincial Quantum Science Strategic Initiative GDZX2404004, and the Space Application System of China Manned Space Program

### v3

@@ Abstract
However, when the model is extended to the quantum realm, the new features arise due to the interplay between quantum coherent time evolution and dissipation. Early investigations into the quantum contact process (QCP) focused on its non-equilibrium phase transition and its distinctions from the classical directed percolation class Marcuzzi2016 ; Buchhold2017 ; Jo2019 ; Roscher2018 ; Carollo2019 ; Gillman2019 ; Gillman2020 ; Jo2021 . A remarkable feature of the QCP is that it can show phase transition in one-dimensional system, however the nature of such a phase transition remains controversial Carollo2019 ; Gillman2019 ; Marcuzzi2016 ; Buchhold2017 ; Jo2019 ; Roscher2018 . Whether the phase transition is continuous or not still remains a matter of debate, and constitutes the main focus of the present work.
@@
Apart from the Liouvillian gap $\mu_{0}$, the structure of the Liouvillian eigenvalues also reflects the time-scale of relaxation towards the steady state. This is quite relevant to the consuming time in a practical numerical simulation. In Fig. 2(b), it is shown the first few largest negative real parts of Liouvillian eigenvalues for $\Omega/\Gamma=5.8$ which is located near the transition point. One can see that the largest negative real part (blue circle) approaches zero as $1/L\rightarrow 0$, as indicated by the arrow Fig. 2(a). Moreover the second largest negative real part (red circles) is constant 0.5-0.5 and it bounds the real parts of all the rest eigenvalues. Namely, the eigenvalues with zero and the largest negative real parts are isolated from the other eigenvalues when the system size becomes larger and larger, signaling a metastability during the time evolution Macieszczak2016 ; Rose2016 ; Macieszczak2021 .
In the inset of Fig. 2(b), it is shown that the time evolution of the averaged population $\langle\bar{n}(t)\rangle$ for $\Omega/\Gamma=5.03$ and $5.035$. Starting with the full active state $|\uparrow\uparrow$$\uparrow\rangle$, the $\langle\bar{n}(t)\rangle$ in both cases decreases fast in the early stage. However it stays in a plateau with finite population for a pretty long time ($10^{2}\Gamma^{-1}$) and then dramatically falls onto the eventual absorbing state. The existence of the metastable state, which is not observed in the (single-site) mean-field approximation, is a consequence of the inclusion of short-range correlations in the system. It reminds us that one should prolong the numerical simulation time to ensure that the system escapes from the metastable state.
@@
Figure 3: (a) The $\langle\hat{n}_{1}\rangle_{\text{ss}}$ as a function of the effective field $F^{n}$ for the CMF simulation with $L=7$. The slope of the thin solid line is one. The intersection of the thin solid line and the curve of $\langle n_{1}\rangle_{\text{ss}}$ denotes the steady-state solution. The intersection denoted by full (empty) magenta diamond indicates the stable (unstable) steady-state solution. (b) The time-evolution trajectories of the reduced state of the boundary site originating from different initial states $|\psi\rangle=\bigotimes_{j}{|\psi_{j}\rangle}$ where the identical $|\psi_{j}\rangles$ locate at the surface of the Bloch sphere. The diamonds and the circle correspond to the intersections in (a). (c) The time evolution of the fidelity between the system state and absorbing state. (d) The time evolution of $\langle\hat{n}_{j}\rangle$ for $\Omega/\Gamma=8$ and $L=11$. The inset of (d) shows the spatial-profile of the steady-state local populations $\langle\hat{n}_{j}\rangle_{\text{ss}}$. (e) The branches of $\langle\bar{n}\rangle_{\text{ss}}$ as a function of $\Omega/\Gamma$ in the CMF approximation with $L=7,9$ and $11$. The upper solid (lower dashed) curves denote the stable (unstable) active states. The red solid line denotes the absorbing state. (f) The resulting CMF transition point $\Omega_{c}(L)/\Gamma$ for different $L$. The green and blue symbols denote the odd- and even-size clusters, respectively. The dashed line indicates the linear fit $\Omega_{c}(L)/\Gamma=-9.61/L+6.04$. The square in magenta marks the transition point $\Omega_{c}/\Gamma=5.83$ through the Liouvillian gap analysis in Fig. 2(a).
@@
Fig. 3(a) shows the steady-state population of the boundary site $\langle\hat{n}_{1}\rangle_{\text{ss}}$ as a function of $F^{n}$ for $L=7$. One can see that for $\Omega/\Gamma=4$ the curve of $\langle\hat{n}\rangle_{\text{ss}}$ only intersects with the straight line $F^{n}=\langle\hat{n}_{1}\rangle_{\text{ss}}$ at the origin meaning that the unique steady state is the absorbing state. While for $\Omega/\Gamma=6$ there are two more intersections at $F^{n}=0.0725$ and $0.3085$ marked by the magenta diamonds. The former (with the derivative greater than one) is an unstable fixed point while the latter indicates the physical steady state. As shown in Fig. 3(b), the unstable steady state is brought to the absorbing state by a random perturbation, while stable steady states tends to attract all the time-evolution trajectories originating from different initial states on the surface of the Bloch sphere. The overlap between the initial state and the absorbing state, which is characterized by the fidelity ${\cal F}=\sqrt{\langle\downarrow\downarrow...\downarrow|\rho_{C}|\downarrow\downarrow...\downarrow\rangle}$, determines which steady state it will evolve. From Fig. 3(c), one can find that the initial states with ${\cal F}>0.89$ are attracted by the absorbing state (${\cal F}\rightarrow 1$ in the long-time limit). To verify whether the CMF effectively recovers the translational invariance, we show the time evolution of the local population $\langle\hat{n}_{j}\rangle$ in the active phase ($\Omega/\Gamma=8$) in Fig. 3(d). One observes that starting from the single-excitation state, the spatial profile of $\langle\hat{n}_{j}\rangle$ becomes homogeneous and the boundary effect is suppressed within the cluster. The stable and unstable steady-state solutions for different $L$ are shown in Fig. 3(e). The feature of the saddle-node bifurcation of the steady-state population is preserved in the CMF results. In particular, the saddle-node point $S_{2}$ remains infinity, while the $S_{1}$ shifts right towards to the transition point in TDL, which evidences the discontinuous phase transition.
The CMF transition points $\Omega_{c}(L)$ for different sizes of cluster is shown in Fig. 3(f). Due to the strong boundary effect in small clusters, there is an even-odd effect in the behavior of the $\Omega_{c}(L)$ as a function of $L$. For odd-size clusters, the $\Omega_{c}(L)$ increases as the $L$ increasing while it exhibits conversely for the even-size clusters. The linear fit of $\Omega_{c}(L)$ as a function of $1/L$ extrapolates the $\Omega_{c}(\infty)\approx 6.04\Gamma$ which agrees well to the transition point $\Omega_{c}/\Gamma\approx 5.83$ predicted by the Liouvillian spectrum analysis LinearFit .
@@ Acknowledgements.
We warmly thank useful discussions with Prof. Fernando Iemini and Dr. Wen-Bin He. This work was supported by Natural Science Foundation of Liaoning Province No. 2025-MS-009 and the Natural Science Foundation of China under Grants No. 11975064. X. L. is supported by the Hong Kong RGC Early Career Scheme (Grant No. 24308323), Guangdong Provincial Quantum Science Strategic Initiative GDZX2404004, and the Space Application System of China Manned Space Program.